-4x-2y=-12
4x+8y=-24

Answer by fractalier(6550)

Okay add these together and get
-4x-2y=-12
4x+8y=-24
--------------
6y = -36
y = -6
now substitute and get
-4x - 2(-6) = -12
-4x + 12 = -12
-4x = -24
x = 6